OpenStudy (anonymous):

If x represents the total number of bags a person checks and each person checks more than two bags, the equation below can be used to determine C, the total fee for all the bags. 20 + 25 + 50(x - 2) = C Arianna paid a total of $145 to check all her bags on a flight. How many bags did she check in all?

OpenStudy (anonymous):

i got 3.8 but i dont think its right ....

OpenStudy (anonymous):

20+25+50x-100=45 45-100+50x=145 -55+50x=145 50x=200 x=4

OpenStudy (anonymous):

ohhhhhhhhhh

OpenStudy (anonymous):

i see what i did wrong